# Proprioceptive State Estimation for Amphibious Tactile Sensing

### v1

@@ 7 Conclusion, Limitations, and Future Works
However, certain limitations deserve consideration. Manufacturing inconsistencies inherent to soft robotics can impact the accuracy of our method. Algorithmic parameters also require precise calibration through physical experiments. Additionally, the rigid plate used for boundary condition acquisition slightly hampers the finger’s compliance, affecting object-finger contact conformity. Finally, the object surface estimation pipeline is sensitive to contact geometry, restricting its use to local surface patches with smooth curvature changes.
@@
Future research seeks to optimize the system for versatile tactile grasping and expand its integration into robotic grippers for diverse land and underwater applications. Furthermore, the vision-based proprioception method holds the potential for developing advanced robotic necks with precise state estimation in amphibious robots driven by parallel mechanisms or pneumatic actuation.
Acknowledgments
This work was partly supported by the Ministry of Science and Technology of China [2022YFB4701200], the National Natural Science Foundation of China [62206119, 52335003], and the Science, Technology, and Innovation Commission of Shenzhen Municipality [JCYJ20220818100417038, ZDSYS20220527171403009].
Appendix
@@ Acknowledgments
  • Movie S1. Evaluation of State Estimation with Motion Capture System. In this movie, we demonstrate four different contact configurations tested using the test platform and present the error measurement protocol utilized in Experiment Section 5.2.
@@
  • Movie S2. Estimating State when Deformed by Touch Haptic Device. In this movie, we showcase the experimental procedures described in Experiment Section 5.3. These involve measuring the position discrepancy between the pen-nib and the nearest node on the soft finger, representing the estimated deformation field error sampled at the corresponding location.
@@
  • Movie S3. Shape Sensing of a Vase Underwater. This movie features a demonstration of the experiment setup described in Experiment Section 5.4. It provides a comprehensive overview of the entire experimental process and showcases the results obtained from estimating the shape of an underwater vase using the soft finger.

### v2

@@ 6 Conclusion, Limitations, and Future Work
However, the study has several limitations. Manufacturing inconsistencies inherent to soft robots can impact the accuracy of our method, and algorithmic parameters require precise calibration through physical experiments. Additionally, using a rigid plate for boundary condition acquisition slightly hampers the finger’s compliance, affecting the contact-based conformation between the object and our finger. The object surface estimation pipeline is also sensitive to contact geometry, restricting its use to local surface patches with smooth curvature changes.
@@
Future research aims to optimize the system for versatile tactile grasping and expand its integration into robotic grippers for diverse on-land and underwater applications. The vision-based proprioception method holds the potential for developing advanced robotic necks for underwater humanoids with precise state estimation driven by parallel mechanisms or pneumatic actuation. These advancements will pave the way for the broader application and utility of vision-based tactile sensing technologies in robotic systems operating in complex environments.
Supplementary Materials
@@
Movie S1. Evaluating Proprioceptive State Estimation using Motion Capture System. This movie demonstrates the four contact configurations tested using the test platform and presents the error measurement protocol utilized in Section 4.1.2.
@@
Movie S2. Estimating State when Deformed by Touch Haptic Device. This movie showcases the experimental procedures described in Section 4.1.3. These involve measuring the position discrepancy between the pen-nib and the nearest node on the soft finger, representing the estimated deformation field error sampled at the corresponding location.
@@
Movie S3. Shape Sensing of a Vase Underwater. This movie features a demonstration of the experiment setup described in Section 4.2.2. It provides a comprehensive overview of the entire experimental process and showcases the results obtained from estimating the shape of an underwater vase using the soft finger.
Movie S4. Vision-based Tactile Grasping with an Underwater ROV This movie demonstrates the experiment in Section 4.2.3, where our soft robotic fingers with in-finger vision installed on the FIFISH Pro V6 Plus robot’s gripper for tactile grasping underwater, providing omni-directional adaptation with real-time finger deformation reconstruction to perceive contact-events underwater.
Appendix A: Abaqus Simulation
The nonlinear deformation computation of the volumetric soft finger is first carried out using Abaqus, an advanced finite element analysis (FEA) software.
Material Calibration
A uniaxial tension test is performed to accurately determine the material’s mechanical properties in the soft finger. The 3rd-order Ogden hyperelastic model is found to be most appropriate for describing the material’s mechanical behavior, as it shows a good match between the experimental stress-strain response and theoretical predictions, as illustrated in Fig. 9.
Figure 9: Modeling hyperelastic behavior of the Hei-cast 8400 using uniaxial tension test data in Abaqus. Several different strain energy potential models, including Mooney-Rivlin, Neo-Hookean, Yeoh, and Ogden (N=3), are selected to fit the test data, and Ogden is the best.
Simulation Setup
In the simulation, one boundary condition, Encastre, secures the finger’s bottom surface. Another boundary condition related to Displacement/Rotation is applied to the node set corresponding to the Aggregation Multi-Handle (AMH) constraints, which deforms the finger. The Abaqus analysis then provides the coordinates of all nodes in the finger mesh.
To enhance the accuracy of simulations depicting the material behavior of Hei-cast 8400, six deviatoric coefficients are incorporated within the Ogden model. However, calibrating these coefficients requires a comprehensive set of mechanical experimental data. The calibration process is further complicated by the variations and inconsistencies typically encountered in the manufacturing processes of soft robotic components [81].
Appendix B: Implementation of the ARAP Method
The As-Rigid-As-Possible (ARAP) method is advantageous in interactive mesh deformation, animation, and 3D modeling, where the objective is to enable users to manipulate shapes while preserving local feature integrity.
Deformation Energy
Utilizing the same notation as in Eq. (6), the element-wise ARAP deformation energy function can be expressed as:
$\Psi_{ARAP}(\Phi_{t_{j}})=\Psi_{ARAP}{(\mathbf{A}_{t_{j}})}=||\mathbf{A}_{t_{j}}-\mathbf{R}||^{2}_{\mathcal{F}},$ (18)
where $\mathbf{R}$ represents the closest rotation to the deformation gradient $\mathbf{A}_{t{j}}\in{\mathbb{R}^{3\times{3}}}$ of the tetrahedron element $t_{j}$ in the Frobenius norm, defined as:
$\mathbf{R}=\mathop{\arg\min}\limits_{\mathbf{R}\in{SO(3)}}||\mathbf{R}-\mathbf{A}_{t_{j}}||^{2}_{\mathcal{F}}.$ (19)
To calculate the deformed positions of the soft finger mesh nodes under the constraints set by AMHs using the ARAP energy model, the constrained geometric optimization problem in Eq. (9) needs to be modified by replacing the symmetric Dirichlet energy Eq. (6) with the ARAP energy in Eq. (18).
Local/Global Optimization
Unlike the symmetric Dirichlet energy discussed in this article, the ARAP energy model does not permit a generic formulation of gradients and Hessians suitable for a Newton-type solver, such as Alg. 1. To adapt the optimization process for deforming soft finger mesh nodes under constraints with the ARAP energy model, we utilized the local-global solver implemented using the libigl library [^3] https://libigl.github.io/, a well-regarded C++ framework known for its efficiency in geometric computations.
The local step concentrates on Eq. (19), aiming to find the closest rotation from the deformation gradient of each mesh element. Subsequently, the global step minimizes the deviation of the deformation gradient from those rotations computed in the local step across all elements. The local and global steps are crucial in maintaining a balance between local shape preservation and the structure’s overall integrity [58].
In addressing the challenge of minimizing deformation energy under AMH constraints, this solver also transforms the constrained problem into an unconstrained one by applying a soft penalty method. We retain the same penalty parameter $\omega$ used in our implementation to ensure a fair comparison.
Appendix C: Parameter Selection for Algorithm 1
As discussed in Section 3.2.4, transforming the constrained geometry optimization problem Eq. (9) into an unconstrained one Eq. (10) by establishing soft boundaries offers several advantages, especially considering the inherent variability in the accuracy of camera-observed constraints. Nonetheless, the potential deviation from constraints due to the soft boundary approach must be meticulously evaluated.
Fig. 10 depicts the relationship between constraint violations and different mesh sizes for a range of penalty parameter $\omega$ values. As $\omega$ increases, there is a consistent reduction in constraint violations, highlighting the effectiveness of adjusting the soft penalty parameter to improve compliance with specified constraints.
Figure 10: Effect of soft penalty parameter $\omega$ on constraint violation.
Moreover, Fig. 11 compares mean errors in geometric optimization against the ground truth established by Abaqus simulations. The data indicates that the mean error experiences less significant fluctuations with larger $\omega$ values.
Based on the results of these empirical evaluations, we also consider the possibility of algorithmic improvements through strategic adjustments to the soft penalty parameter $\omega$ strategy. Specifically, we suggest that adopting an adaptive scheme, where $\omega$ is systematically increased with each iteration, or employing the augmented Lagrangian method to dynamically optimize $\omega$ for each iteration, could significantly enhance the performance of our algorithm. These approaches are expected to enable more accurate constraint handling and improve overall shape estimation accuracy by ensuring the optimal balance of $\omega$ in response to the evolving conditions of the optimization process.
For this study, we have utilized a constant $\omega=10^{5}$, motivated by our aim to establish a stable baseline to evaluate the core capabilities of our method. This fixed value was chosen to provide consistency across our experiments, facilitating a more straightforward assessment of the foundational performance of our shape estimation method.
Figure 11: Effect of soft penalty parameter $\omega$ on shape reconstruction accuracy.